Opened 8 years ago

Last modified 6 months ago

#283 reopened flight processing

ET07/01, flight day 299/2009, Afar north

Reported by: benj Owned by: knpa
Priority: alpha 5 Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/ethiopia/ET07_01-2009_299_Afar_north

Data arrived from ARSF via SATA disk Y on 4th Nov. 2009.

Scientific objective: /users/rsg/arsf/arsf_data/2008/ARSF_Applications/Ethiopia/ET07_01\ -\ Dr\ Wright.pdf

Priority: alpha-5

PI: Tim Wright

2009 Ethiopia campaign, reflown from 2007

Sensors:

  • Eagle (23/04/10)
  • Hawk (23/04/10)
  • Leica LIDAR (19/04/10)
  • Photography (23/04/10)
  • TABI (to be processed by University of Athens)

Change History (27)

comment:1 Changed 8 years ago by benj

  • Description modified (diff)

comment:2 Changed 8 years ago by benj

Raw GPS base station data is in ~arsf/arsf_data/2009/flight_data/ethiopia/basestations.

comment:3 Changed 8 years ago by mggr

  • Description modified (diff)

Raw TABI data acquired, put in archive and offered to Athens for download and processing.

comment:4 Changed 8 years ago by knpa

  • Owner set to knpa
  • Status changed from new to accepted

comment:5 Changed 8 years ago by mark1

No GPS time in event log file for photograph images:

25222049100005G3.raw
25222049100006G3.raw

Times have been estimated from SensorStat file

comment:6 Changed 8 years ago by mark1

  • Description modified (diff)

comment:7 Changed 8 years ago by knpa

LIDAR processed and QCd.
Needs to be classified and can then be delivered.
Passing on to harg to continue.

comment:8 Changed 8 years ago by knpa

  • Owner changed from knpa to harg
  • Status changed from accepted to assigned

comment:9 Changed 8 years ago by knpa

Pitch offset changed during QC to -0.00058

comment:10 Changed 7 years ago by mark1

Digital photography is ready to be delivered

comment:11 Changed 7 years ago by chrfi

hawk data for flight line 1 is not present on thelma or our systems.

comment:12 Changed 7 years ago by emca

LIDAR data processing complete, delivery directory created at
~airborne/workspace/ET07_01-2009_299_Afar_north/delivery/20100415/ET07_01

Delivery checked by mark1 16/04/2010

Data dispatched to PI 19/04/2010

comment:13 Changed 7 years ago by knpa

Hyperspectral Delivery Checked

All is fine apart from following:

Need to mention eagle dropped scan lines in Readme
Need to mention clouds on a couple of flightlines (e.g. h24, e16, e18, e22, e23, e24, e25)
Can't open eagle mosaic jpeg (with eog)
bin dir not mentioned in contents
wonky | on line 227
In contents should say line name, not number e.g. Pol1, not 01.
"indicatethe accuracy" - line 277
need to unix2dox Readme
Need to mention blank 232 band in all hawk flight lines

Ready to deliver when these changes are made.

comment:14 Changed 7 years ago by chrfi

hyperspectral time offsets

line number eagle hawk
1 0.06 n/a
2 -0.02 -0.04
3 0.07 -0.02
4 -0.02 0.02
5 0.00 0.01
6 -0.05 -0.05
7 -0.01 0.00
8 0.01 -0.02
9 -0.02 -0.06
10 -0.05 -0.08
11 0.03 0.00
12 0.02 -0.02
13 0.00 -0.04
14 0.00 -0.04
15 0.00 0.00
16 0.02 -0.02
17 0.00 0.00
18 -0.01 -0.06
19 0.00 -0.04
20 0.35 0.7
21 -0.02 0.02
22 -0.03 -0.02
23 0.04 -0.01
24 -0.03 -0.01
25 0.00 -0.04

comment:15 Changed 7 years ago by chrfi

dispatched 04/23/2010

comment:16 Changed 7 years ago by benj

Processed TABI data was received from Pol back in January but we weren't happy with the accuracy. This data is in the workspace, it has not been dispatched to the PI but is the only processed TABI data we currently have for this project.

comment:17 Changed 7 years ago by knpa

  • Description modified (diff)

Photography data too!
Also added delivery dates from processing spreadsheet.

comment:18 Changed 7 years ago by knpa

  • Owner changed from harg to knpa

r-synced all sensors.

Beginning archiving.

comment:19 Changed 7 years ago by knpa

Found raw data for hawk line 1, it was misnamed 'SWIRdataCube' (there is already the normal dataCube.raw present).

Holding archiving whilst I process and stick on ftp for PI.

comment:20 Changed 7 years ago by knpa

fps has already been changed from 60 to 30 for all hawk header files due to to frame dropping issue. Have changed for line 1 also.

comment:21 Changed 7 years ago by knpa

hawk line 1 sct: -0.01

comment:22 Changed 7 years ago by knpa

Made line available to PI on ftp 20/08/10

comment:23 Changed 7 years ago by knpa

Archiving preparation complete.

Tarball created and ready for download by NEODC.

comment:24 Changed 7 years ago by knpa

Tarball has been downloaded and the data backed-up by NEODC.

Project has been moved from repository to archive disk.

comment:25 Changed 7 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:26 Changed 6 years ago by mark1

  • Component changed from Processing: general to Archiving
  • Resolution fixed deleted
  • Status changed from closed to reopened

Eagle and Hawk line 20 re-delivered in both az and apl formats. (Put up on FTP 1st July 2011). Ticket re-opened as these will need to be archived.

comment:27 Changed 6 months ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.